OpenStudy (anonymous):

i need to know how adding and subtracting rational expressions is similar to adding and subtracting simple fractions and how they are different

OpenStudy (isaiah.feynman):

Ponder on this simple formula..\[\frac{ a }{ b }\pm \frac{ c }{ d }= \frac{ ad \pm bc }{ bd }\] THIS FORMULA IS UNIVERSAL, IT APPLIES TO BOTH RATIONAL EXPRESSIONS AND SIMPLE NUMERICAL FRACTIONS.

OpenStudy (anonymous):

im not sure how to put that in words

OpenStudy (anonymous):

im not even sure what that is

OpenStudy (isaiah.feynman):

IT SHOWS HOW TO ADD OR SUBTRACT TWO FRACTIONS.

OpenStudy (anonymous):

so you add the two fractions and then you put them both together?

OpenStudy (isaiah.feynman):

A quick example with simple fractions..